OpenStudy (anonymous):

16sqrt((-8))^(16) simplify

OpenStudy (anonymous):

\[16\sqrt{-8}^{16}\]?

OpenStudy (anonymous):

\[16(\sqrt{-8})^{16}\]?

OpenStudy (anonymous):

\[(\sqrt{(-8)})^2 = +8, so (\sqrt{(-8)})^(16)= 8^8, so answer is 16 \times 8^8, or2\times 8^9\]

OpenStudy (anonymous):

That's 8 to power 9